OpenStudy (anonymous):

solve [y=x^2-3x+6 y=-x^2+7X+6

OpenStudy (n00bstyle):

You can see you have two equations with two unknowns. This means you can solve this (because you have the same number of equations, as you have unknows). Both are expressed in y. The only thing you have to do is substitute the two equation, can you do that?

OpenStudy (n00bstyle):

For example, when you have y=3 and also a formula y=4+x This means you will have to calculate: 3=4+x Now you can calculate x. I now substitued y=3 in the second equation

OpenStudy (mathmate):

Another way is to solve it by comparison: Given y=x^2-3x+6 y=-x^2+7X+6 Since the left-hand sides are both equal to y, we can equate the right-hand sides: x^2-3x+6 = -x^2+7x+6 All we have to do now is to simplify the above to 2x^2-10x = 0 or x^2-5x =0 Now solve for x (2 solutions) and substitute x in the original equations to find y.